Credit Card Sales Historically, 60 percent of the customer bills at the Andrews’ Supper Club have been paid with cash or check, and 40 percent have been paid using either the Great American Bank Card or the United Merchants Card. Andrews pays a four percent fee with both cards. Great American Bank deposits cash in Andrews’ checking account when the credit card sales slips are deposited. United Merchants makes an electronic funds transfer three days after the sales slips are mailed. Prepare journal entries to record the following: Round to the nearest dollar.

a. Sales for September 10 were as follows: Cash and checks $1,340 Great American Bank Card (Deposited at the end of the day) 500 United Merchants Card (Mailed at the end of the day) 300 $2,140

b. On September 13, Andrews received an electronic funds transfer from United Merchants for the September 10 sales.
Answer #1

Answers @ sept-10: cash = 1340 +(500 x 964) = 1340 +(500X0.96) = 1340 X 480 = 1,820 (3 60. + 404 –44 ) =964 Accounts Receivabsept to 1920 288 - Cash accounts Receivcible - united merchards Credit card tee Expense sales Revenue 2140 sept13 288 Cash 28